# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v3

@@ Appendix A Additional Tables
 [34] III brady brady from of gait stability posture tremor tremor tremor tremor tremor
 [35] kinesia kinesia chair composite components
in-air duration **** ** n.s. n.s. n.s. n.s. ** * n.s. ** n.s. n.s. n.s. * ** n.s.
in-air avg. X speed *** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** *** **** n.s.
in-air std X speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * **** **** n.s.
in-air amp X speed * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** *** n.s.
in-air avg. X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** *** *** n.s.
in-air std X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** *** **** n.s.
in-air amp X acc. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * * n.s.
in-air avg. Y speed *** n.s. n.s. n.s. n.s. * * n.s. n.s. n.s. n.s. * ** *** *** n.s.
in-air std Y speed n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. **** *** n.s.
in-air amp Y speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** *** n.s.
in-air avg. Y acc. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** *** *** n.s.
in-air std Y acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * *** **** n.s.
in-air amp Y acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** ** n.s.
in-air avg. speed ** n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** * *** *** n.s.
in-air std speed ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** * n.s.
in-air amp speed *** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** ** n.s.
in-air avg. acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** * *** *** n.s.
in-air std acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** * n.s.
in-air amp acc. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** ** n.s.
in-air std angular speed **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. * n.s.
in-air avg. angular speed **** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. * * n.s. * n.s.
in-air std angular acc. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s.
in-air avg. angular acc. *** n.s. * n.s. n.s. n.s. ** n.s. n.s. n.s. n.s. * * n.s. n.s. n.s.
in-air shannon entropy **** n.s. n.s. n.s. n.s. * * * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet duration **** ** n.s. n.s. n.s. n.s. ** * n.s. n.s. n.s. n.s. n.s. * ** n.s.
on-tablet avg. X speed ** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std X speed * n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp X speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. X acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp X acc. *** n.s. n.s. n.s. n.s. n.s. n.s. **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. Y speed * n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std Y speed * n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp Y speed n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. Y acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std Y acc. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp Y acc. ** n.s. n.s. n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. speed n.s. n.s. * n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std speed n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp speed n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std acc. n.s. * n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp acc. n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std angular speed **** n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
on-tablet avg. angular speed **** ** * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std angular acc. **** ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** ** n.s.
on-tablet avg. angular acc. *** ** n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. ** n.s. * n.s.
on-tablet avg. dP/dt ** n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s.
on-tablet std dP/dt * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** ** n.s.
on-tablet amp dP/dt * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet avg. d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet std d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** ** n.s.
on-tablet amp d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet shannon entropy **** n.s. n.s. n.s. n.s. n.s. ** **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
TABLE XII:

### v4

@@ Appendix A Additional Tables
name bradykinesia bradykinesia from chair of gait stability posture composite tremor tremor tremor tremor components
in-air duration **** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s.
in-air avg. X speed * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** n.s. * n.s.
in-air std X speed * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. * n.s.
in-air amp X speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
in-air avg. X acc. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. * n.s.
in-air std X acc. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
in-air amp X acc. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
in-air avg. Y speed * n.s. n.s. n.s. n.s. ** * n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
in-air std Y speed n.s. n.s. n.s. n.s. n.s. ** * n.s. n.s. n.s. n.s. n.s. n.s. n.s. * *
in-air amp Y speed * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
in-air avg. Y acc. * n.s. n.s. n.s. n.s. *** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
in-air std Y acc. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
in-air amp Y acc. ** n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
in-air avg. speed **** n.s. n.s. n.s. n.s. ** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. * n.s.
in-air std speed **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. * n.s.
in-air amp speed **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. * n.s.
in-air avg. acc. **** n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. * n.s. n.s. * n.s.
in-air std acc. **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. * n.s.
in-air amp acc. **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
in-air std angular speed ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
in-air avg. angular speed ** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. * * n.s. * n.s.
in-air std angular acc. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
in-air avg. angular acc. ** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. * * n.s. n.s. n.s.
in-air shannon entropy **** n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
on-tablet duration *** n.s. n.s. ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s.
on-tablet avg. X speed ** n.s. n.s. n.s. * * * n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
on-tablet std X speed ** n.s. n.s. n.s. * * n.s. n.s. * n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp X speed n.s. n.s. n.s. n.s. ** n.s. n.s. ** ** n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet avg. X acc. n.s. n.s. n.s. n.s. * * * n.s. n.s. n.s. n.s. * n.s. n.s. * n.s.
on-tablet std X acc. n.s. n.s. n.s. n.s. * n.s. n.s. ** * n.s. n.s. * n.s. n.s. * n.s.
on-tablet amp X acc. n.s. n.s. n.s. n.s. * n.s. n.s. ** n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet avg. Y speed n.s. n.s. n.s. n.s. * * * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std Y speed * n.s. n.s. n.s. * * * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp Y speed n.s. n.s. n.s. n.s. * n.s. * * * n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. Y acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. * n.s.
on-tablet std Y acc. n.s. n.s. n.s. n.s. ** n.s. n.s. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp Y acc. n.s. n.s. n.s. n.s. ** n.s. n.s. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. speed n.s. n.s. n.s. n.s. *** * * n.s. ** n.s. n.s. * n.s. n.s. * n.s.
on-tablet std speed n.s. n.s. n.s. n.s. ** n.s. * * n.s. n.s. n.s. n.s. n.s. * * n.s.
on-tablet amp speed * n.s. n.s. n.s. ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
on-tablet avg. acc. n.s. n.s. n.s. n.s. *** n.s. * * ** n.s. n.s. * n.s. n.s. * n.s.
on-tablet std acc. n.s. n.s. n.s. n.s. ** n.s. n.s. * n.s. n.s. n.s. * n.s. * * n.s.
on-tablet amp acc. n.s. n.s. n.s. n.s. ** n.s. n.s. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std angular speed * n.s. n.s. n.s. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. angular speed **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std angular acc. n.s. n.s. n.s. n.s. ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. angular acc. n.s. n.s. n.s. n.s. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. * n.s.
on-tablet avg. dP/dt n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. * ** n.s.
on-tablet std dP/dt n.s. n.s. n.s. n.s. ** n.s. * n.s. n.s. n.s. n.s. * n.s. * * n.s.
on-tablet amp dP/dt n.s. n.s. n.s. n.s. ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. d2P/dt2 n.s. n.s. n.s. n.s. ** n.s. * n.s. n.s. n.s. n.s. * n.s. * * n.s.
on-tablet std d2P/dt2 n.s. n.s. n.s. n.s. ** n.s. * n.s. n.s. n.s. n.s. * n.s. * * n.s.
on-tablet amp d2P/dt2 n.s. n.s. n.s. n.s. ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet shannon entropy *** n.s. n.s. ** * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
TABLE XII: Table of classification performances for the Random Forest and Multi Layer Perceptron classifiers.